How do we expect the luxury package to affect price? Round to nearest whole number The dependent variable is the sales price of a used car. The independent predictor variables are and whether or not the car has a luxury package (yes 1, mileage 0). Answer questions 31 -32 no Regression Stetistics mileage luxury package salesprice 42,000 Multiple R 0.90 26,500 0.31 R 5qure 2 37,500 29,400 Adjusted R Square 0,78 41430 24,000 33,700 Standard Emor 3577.08 19,670 1 ODEervations 15 53,800 48 900 18,550 C 27,200 ANOVA 28,750 35,400 MS 14.220 38.550 s89144028.3 2.95E+08 25.8198504 Regression 12 880 33,000 Residual 12 15004905 114087421 10 64,330 39.890 14,990 Tota 14 7260433333 11 0 16,770 12 40,100 28,750 Coefficients Standard Eror P-value t Stat 13 25,500 32,990 intercept 3847.34 2465.65 15.80 Q.00 14 15.320 35,900 mileage 0.36 0.06 -6.04 0.00 15 17,860 31,550 1s0. luxury package s557.91 2.10 005 The average sales price is expected to increase by about $3,888 The average sales price is expected to increase by about $1,851 It has no impact on price by itself It explains 819% of the price None of the above
A quality control expert sampled 60 parts from a production line. If the part has a defect, the defect is classified as either minor or major. Of the 60 sampled parts, 4 had minor defects and 2 had major defects. Use this information for questions 33-34. What is the probability that a part will have some type of defect (any type? 8% 10% 16.6% 14.3% None of the above
A quality control expert sampled 60 parts from a production line. If the part has a defect, the defect is classified as either minor or major. Of the 60 sampled parts, 4 had minor defects and 2 had major defects. Use this information for questions 33-34 Given that a part has a defect, what is the probability that the defect will be minor (round to the nearest decimal)? 62.5% 66.7% 33.3% 50% None of the above
Answer #1

Answer:

The average sales price is expected to increase by about $3888.

(Regression coefficient for luxury package is 3887.91)

Problem 33

P( some type of defect)= 6/60 = 0.10   or

10%

Problem 34

P( minor defect/ has a defect)= 4/6 = 0.667 or

66.7%